OpenStudy (anonymous):

let f(x)=x^2+5 and G(x)= (x+5) / x find (g degree f) (-3)

OpenStudy (anonymous):

g degree f means g composition f?

OpenStudy (anonymous):

f(-3) = 9+5 = 14 (g degree f) = g (14) = (14+5)/ 14 = 19/5 = 3.8